OpenStudy (anonymous):

Algebra 2 Help, Please! (cos Θ + cos Θ)2 + (cos Θ + cos Θ)2

OpenStudy (michele_laino):

hint: if we compute the summations, we get: \[{\left( {2\cos \theta } \right)^2} + {\left( {2\cos \theta } \right)^2} = ...?\]